An antique dealer found a 14K gold statue at an auction. In order to maximize profits from this statue, it was melted down with cheap 6K gold trinkets to produce 8K gold replica. If the replica has a mass of 3.2Kg, what were the masses of the statue and the trinkets used?
Answer by checkley71(8403)

14x+6(x-3.2)=8*3.2
14x+6x-19.2=25.6
20x=25.6+19.2
20x=44.8
x=44.8/20
x=2.24 kg was the weight of the statue.
3.2-2.24=.96 kg is the amount of the 6K gold used.
proof:
14*2.24+6*2.24-19.2=25.6
31.36+13.44-19.2=25.6
25.6=25.6